OpenStudy (torvela):

Solve for x: −4|x + 5| = −16

OpenStudy (torvela):

@Mehek14

Mehek (mehek14):

First we have to get rid of the -4 Divide both sides by -4

OpenStudy (torvela):

so i ended up with x=17 and x=7 but thats not an answer choice

OpenStudy (torvela):

wait wouldn't you just add 4?

Mehek (mehek14):

Nope

OpenStudy (torvela):

ok thats where i messed up

Mehek (mehek14):

-4 is being multiplied to |x+5| so you have to divide in order to move it to the other side

OpenStudy (torvela):

ok now I'm at |x-5|=-4

OpenStudy (torvela):

is that right so far

Mehek (mehek14):

It should be |x+5| = 4 -16/-4 = 4

OpenStudy (torvela):

ok

Mehek (mehek14):

So now we need to equations first one being x + 5 = 4 and the second one x + 5 = -4

Mehek (mehek14):

Solve for x in both

OpenStudy (torvela):

ok

OpenStudy (torvela):

so do i start by subtracting each side by 5

OpenStudy (torvela):

because this doesn't seem quite right

OpenStudy (torvela):

Mehek?

Mehek (mehek14):

It is right

Mehek (mehek14):

Since it's being added on one side, you subtract so you can move it to the other

OpenStudy (torvela):

what i don't understand what you mean

Mehek (mehek14):

Alright so you see how it is x + 5?

Mehek (mehek14):

We want to get x by itself so you subtract -5 on both sides

Mehek (mehek14):

This way we eliminate the +5 x + 5 = 4 x + 5 - 5 = 4 - 5 x = 4 - 5

OpenStudy (torvela):

so x is equal to -9 and -1?

Mehek (mehek14):

Yep

OpenStudy (torvela):

thanks so much!

Mehek (mehek14):

Plug them into the original equation to see if they're solutions

Mehek (mehek14):

No problem (: